# A study on the adequacy of common IQA measures for medical images

### v2

@@ 4 Conclusion
We systematically test the ability of common FR and NR IQA measures to assess the quality of medical images and find that all measures correlate less with expert assessment of medical images compared to grayscale natural images. Even though PSNR and SSIM are commonly used across fields, they yield relatively low results, confirming previous studies that these two measures are not beneficial to assess medical imaging tasks and emphasizing the need to employ specifically task-targeted IQ measures. On the other hand, HaarPSI showed exceptional behavior regarding generalizability, suggesting it acts as a robust measure of quality in addition to such specifically tailored measures. FSIM, GMSD and MS-SSIM also showed relatively robust successful behavior across the data sets; LPIPS and DISTS succeeded in most tasks.
In summary, we present a comprehensive study to assess the adequacy of common IQA measures in medical imaging tasks including rated chest X-ray scans and photoacoustic images. We show the dire need for better suited IQA measures and hope to provoke more research in this direction, including more extensive suitability studies as well as further implementations of task-tailored measures.

### v4

@@ 5 Conclusion
We systematically test the ability of common FR and NR IQA measures to assess the quality of medical images and observe that the measures correlate less with expert assessment of the medical images compared to the grayscale natural images. Even though PSNR and SSIM are commonly used across fields, they yield relatively low results. This confirms previous studies, showing that these two measures are not necessarily a beneficial choice to assess medical imaging tasks, and emphasizing the need to employ task-tailored IQ measures more often. On the other hand, HaarPSI showed exceptional behavior regarding generalizability, suggesting that it may act as a robust measure of quality across domains to complement task-tailored measures. FSIM, GMSD and MS-SSIM also showed relatively robust successful behavior across the data sets; LPIPS and IW-SSIM succeeded in most tasks.
In summary, we presented a comprehensive study to assess the adequacy of common IQA measures in medical imaging tasks including rated chest X-ray scans and photoacoustic images, which are made publicly available. We show the dire need for better suited IQA measures and hope to provoke more research in this direction, including extensive suitability studies, adaption of existing measures for medical images as well as further implementation of task-tailored measures.